The demand for a product is unit elastic. At a price of $20, 10 units of a product are sold. What would one expect sales to equal if the price is increased to $40? Be sure to provide an explanation


1
Expert's answer
2022-03-01T14:54:09-0500

It is given that demand is unit elastic. It means it is equal to 1. It is given that the pries of $20, 10 units are sold

Now. price increased to $40.

When the demand is unit elastic that is 1, the total revenue remains the same even if the price increases or decrease

Initial total revenue.

TR ="\$20\times 10= \$200"

It means that the total revenue is $200 when the price increases to $40.

200 = "\$40\times Q"

"Q=\frac{200} {40}" = 5